Eighteen years ago, Porshia’s grandmother bought her a $1,000 face value bond for $500. Today, she can cash this bond in for $1,514.67. What is the effective interest rate earned by this bond?
Answer #1

Ans 6.35%

Eighteen years ago, Porshia’s grandmother bought her a $1,000 face value bond for $500. Today, she can cash this bond in for $1,514.67. The effective interest rate earned by this bond is 6.35%

FV = Future Value PV = Present Value r= rate of interest n= no of period

FV/ PV = (1+r)^n 1514.67/ 500 = (1 +r%)^18 3.02934 = (1 +r%)^18 r= 6.35%